OpenStudy (error1603):

Wallace ran 100 meters in 12 seconds at a track meet. What was Wallace's speed, in Kilometers per hour?

OpenStudy (error1603):

1000 meters = 1 kilometer 3600 seconds = 1 hour So, 100 meters = 0.1 kilometer 12 seconds = 1/300 hour 0.1 kilometer = 1/300 hour x = 1 hour x/300 = 0.1 x=30 kilometers per hour Wallace ran 30 kilometers per hour

OpenStudy (anonymous):

Yeah